OpenStudy (anonymous):

Solve The Equations Below ; 4=7+x+6x -3/4m+3=8

OpenStudy (anonymous):

x=-3/7 first 1

OpenStudy (anonymous):

Thankss (:

OpenStudy (anonymous):

second 1 = -20/3

OpenStudy (anonymous):

understan why?

OpenStudy (anonymous):

Noo.

OpenStudy (anonymous):

hang on ill explain

OpenStudy (anonymous):

Alrightt '

OpenStudy (anonymous):

\[4=7+x+6x\] \[4=7+7x\] combine like terms \[-3=7x\] minus 7 from both sides \[x=-3/7\] get the x by itsel by deviding by 7

OpenStudy (anonymous):

sorry for delay im new at typing on this thing

OpenStudy (anonymous):

the second equation is the same process

OpenStudy (anonymous):

Hahah ' Thanks For The Helpp (: